OpenStudy (anonymous):

Use Pascal’s triangle to find the number of ways to choose 5 shirts from 7 shirts. a. 6 c. 21 b. 56 d. none of these

OpenStudy (anonymous):

the bottom row on the second triangle is the 7th level http://www.mathsisfun.com/pascals-triangle.html

OpenStudy (anonymous):

and your answer is highlighted in dark blue

OpenStudy (anonymous):

This can be verified by \[\left(\begin{matrix} 7\\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

Which is the same as \[\frac{ 7! }{ 5! \times (7-5)! } = 21\]